# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v2

@@ 4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin et al., 2024). Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy. In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt. The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
@@ 4.1. Experiment Setup
To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples. As described in section 3, we design a four-stage data cleaning pipeline. Stage 0 includes basic textual cleaning. Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar, 2024). Stage 2 excludes short reviews with three words or less. Stage 3 excludes non-informative reviews using an LLM-based filtering approach. We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 4 and detail it in this section.
@@
Topic extraction. To evaluate how prompt design affects topic extraction performance, we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos Santos et al., 2023; Assi et al., 2025a; Pham et al., 2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instruction with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. Figure 5 shows the three prompt configuration templates.
@@
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the set of top topics $T(x,y,n)$ extracted in that experiment, 2) five few-shot example reviews manually selected and labelled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Assi et al., 2025a; Deshmukh et al., 2025).
@@
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (Cohen, 1960) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labelling criteria. In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan and Walmsley, 1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann and Whitney, 1947) with Bonferroni correction to adjust for multiple testing (Cabin and Mitchell, 2000). We also use one-way ANOVA (Judd et al., 2017) to test for significant differences across cleaning stages and few-shot settings.
4.2. RQ1 Findings
Table 3. Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video).
@@ 4.2. RQ1 Findings
Table 4. Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video).
@@
5. RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps. In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories. We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.
@@ 5.1. Experiment Setup
Topic extraction and assignment. The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration. Our dataset spans 11 app categories, totalling 676,066 user reviews as shown in Table 2. As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error. The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category. For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5. Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset. For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
Classification and categorization. As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features. To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI. We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns. In line with prior studies (Assi et al., 2025a; Pagano and Maalej, 2013; Chen et al., 2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
Analysis. We perform quantitative analysis on both app categories and topic categories. We compute the following metrics:
@@
    Average Gen-AI Rating (AGR): the average star rating of reviews labelled with Gen-AI topics.
@@
    Average Non-Gen-AI Rating (ANR): the average rating of reviews labelled with non-Gen-AI topics.
@@
    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI
@@
    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics
@@
    Gen-AI Review Percentage (GRP): the proportion of each type relative to the total number of reviews in each app category
@@
    Non-Gen-AI Review Percentage (NRP): the proportion of each type relative to the total number of reviews in each topic category
We conduct the Mann–Whitney U test to examine whether reviews labelled as Gen-AI-related receive significantly higher ratings than those that are not.
To gain deeper insights into user concerns and feedback within each topic category, we randomly sample reviews using a 95% confidence level and a 10% margin of error, resulting in a dataset of 901 reviews. We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
@@ 5.2. RQ2 Findings
Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories. As shown in Table 5, the average rating of reviews discussing Gen-AI functionality exceeds the average rating of non-Gen-AI reviews in every app category, indicating higher user satisfaction with the Gen-AI features.
To assess the statistical significance of this difference, we apply the Mann–Whitney U test. The results reveal a significant difference in the rating distributions between the two groups ($U=65{,}113{,}513{,}850$, $p<0.0001$). Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews. These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
Table 5. Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics
@@
Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category. As reported in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains. For example, categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively. Conversely, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews. In these categories, 58% of the reviews focus on non-Gen-AI topics, such as overall user experience, technical issues, and subscription or pricing models, suggesting that these factors outweighed interest in Gen-AI features.
Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting key differences in engagement. Several categories, such as Education and Productivity, exhibit a moderate spread, indicating consistent Gen-AI engagement across apps. Other categories, such as Art & Design and Entertainment present a wider spread, suggesting more variability in Gen-AI adoption among apps within these groups.
@@
Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category
$\Rightarrow$ Implication 1.
Gen-AI reviews consistently yield higher user ratings across app categories. Developers must recognize Gen-AI as a crucial tool for boosting overall user satisfaction. They must strategically incorporate Gen-AI features to remain competitive.
@@
Observation 2.3: The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality. Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively. In terms of user satisfaction, the Gen-AI topics with the highest average ratings are Human AI Interaction & Emotional Connection, Utility & Use Cases, and Customer Support & Community. The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship. We further examine each of these categories in the subsections below.
@@
Table 6. Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears
@@
Table 7. Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears
@@
$\Rightarrow$ Implication 2.
While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts. To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions.
@@
However, 33% of reviews reveal concerns in three domains. First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk et al., 2022). Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion. For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour. Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
$\Rightarrow$ Implication 3.
Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content. Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
@@
$\Rightarrow$ Implication 5.
Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them. For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences.
@@
$\Rightarrow$ Implication 6.
Despite the launch of tools like ChatGPT Edu (OpenAI, 2024a), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features. Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
@@
$\Rightarrow$ Implication 7.
Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows. They should also add inline editing tools, so users can refine content directly without needing to regenerate responses. Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized.
@@
$\Rightarrow$ Implication 8.
ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers. Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search. The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary).

### v3

@@ 4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; 10.1145/3736408). Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy. In our setup, we adopt this principle by refining prompts through multiple pilot runs on statistically significant samples to improve label consistency and clarity. In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt. The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
@@ 4.1. Experiment Setup
To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples. As described in section 3, we design a four-stage data cleaning framework. Stage 0 includes basic textual cleaning. Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar_2024). Stage 2 excludes short reviews with three words or less. Stage 3 excludes non-informative reviews using an LLM-based filtering approach. We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 4 and detail it in this section.
@@
Topic extraction. Because prior work reported mixed effects of few-shot examples (Wang_Yang_Wei_2023; Prakash_Wang_Hoang_Hee_Lee_2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023; Maram_LLM_Cure; Pham_Hoyle_Sun_Resnik_Iyyer_2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instructions with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. Figure 5 shows the three prompt configuration templates.
To mitigate hallucination during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention. To further adapt the LLM to the Gen-AI domain, the few-shot examples are drawn from real reviews of Gen-AI apps, exposing the model to characteristic terminology and contexts (e.g., discussions of AI Performance, Technical Difficulties, and Content Quality).
@@
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Maram_LLM_Cure; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025).
To mitigate hallucination during topic assignment, we constrain the model’s output to a predefined set of deterministic labels, i.e., the extracted topics $T(x,y,n)$. The model must choose from these labels or select “Other” if none accurately represent the review. This closed-set format enables immediate detection of hallucination, as any label outside the predefined list signals a deviation from valid outputs. Additionally, structured CSV-formatted output further enforces adherence to the expected schema and prevents free-form or unconstrained generation. To ensure domain relevance, the five few-shot examples are handpicked from real reviews of Gen-AI apps, representing common contexts and linguistic patterns found in this domain (e.g., user reviews of AI Performance, Utility & Use Cases, or Content Quality).
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (cohen1960kappa) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labeling criteria. In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan_Walmsley_1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann_Whitney_1947) with Bonferroni correction to adjust for multiple testing (3b6386c9-eeae-33a7-8c59-c0604aa87bef). We also use one-way ANOVA (Judd_McClelland_Ryan_2017) to test for significant differences across cleaning stages and few-shot settings.
Table 3. Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video). (Representative subset considered).
@@ 4.2. RQ1 Findings
Table 4. Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video). (Representative subset considered).
@@
Observation 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified. Among these, 2% of all reviews involved incorrect topic assignments, 3% were missed topics (valid reviews marked as “Other”), and 4% were over-assignments (non-topical reviews given a label). Most errors occurred in short or ambiguous reviews with limited context, where topic boundaries were less clear. These findings indicate that misclassifications are infrequent and stem from understandable edge cases rather than uncontrolled generation or hallucination, reinforcing the reliability of the LLM-based labeling process.
@@
5. RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
Having validated the accuracy of our LLM-based topic extraction and assignment framework in RQ1, we now apply it to the full dataset to examine what users discuss in their reviews of Gen-AI apps. Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps. In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories. We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.
@@ 5.1. Experiment Setup
Topic extraction and assignment. The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration. Our dataset spans 11 app categories, totaling 676,066 user reviews as shown in Table 2. As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error. The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category. For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5. Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset. For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
Classification and categorization. As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features. To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI. All reviews originate from Gen-AI apps; the distinction here is between review types, not between Gen-AI and non-Gen-AI apps. We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns. In line with prior studies (Maram_LLM_Cure; Pagano_Maalej_2013; Chen_Lin_Hoi_Xiao_Zhang_2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
Analysis. We perform large-scale quantitative analysis on all automatically labeled reviews across both app categories and topic categories, followed by qualitative analysis on a representative sample. Specifically, for the quantitative component, we compute the following metrics:
@@
    Average Gen-AI Rating (AGR): the average star rating of reviews labeled with Gen-AI topics.
@@
    Average Non-Gen-AI Rating (ANR): the average rating of reviews labeled with non-Gen-AI topics.
@@
    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI topics.
@@
    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics.
@@
    Gen-AI Review Percentage (GRP): the proportion of Gen-AI reviews relative to the total number of reviews.
@@
    Non-Gen-AI Review Percentage (NRP): the proportion of non-Gen-AI reviews relative to the total number of reviews.
We conduct the Mann–Whitney U test to examine whether reviews labeled as Gen-AI-related receive significantly higher ratings than those that are not. For developer-response analysis, we compute reply ratios, response delays, and response-type distributions for both Gen-AI and non–Gen-AI reviews, and we compare per-app reply behavior using a Wilcoxon signed-rank test (wilcoxon1992).
We complement our quantitative analyses with qualitative examination by manually coding statistically significant samples (95% confidence level, 10% margin of error) of both user reviews and developer responses. For user reviews we draw samples from every Gen-AI topic category resulting in a total of 901 reviews. We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
For developer responses, we perform two qualitative analyses. First, we code a representative sample of 96 replies drawn across all app categories to characterize the overall distribution of developer responses. Second, we manually code an additional 281 replies from three focal categories (Games, Education, and Entertainment) to capture deeper, category-specific response practices. Together, these qualitative analyses provide contextual grounding for the quantitative findings and offer a more complete understanding of both user concerns and developer engagement within Gen-AI apps.
@@ 5.2. RQ2 Findings
Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories of Gen-AI apps. As shown in Table 5, the average rating of reviews discussing Gen-AI functionality (i.e., AI Performance and Utility & Use Cases) exceeds the average rating of non-Gen-AI reviews (i.e., Technical Difficulties and Monetization Methods & Structure) in every app category of Gen-AI apps. This indicates that, when users explicitly comment on Gen-AI features, they tend to rate their experiences more positively than when they comment on other aspects of the same apps.
To assess the statistical significance of this difference, we apply the Mann–Whitney U test. The results reveal a significant difference in the rating distributions between the two groups (U=65,113,513,850U=65{,}113{,}513{,}850, $p<0.0001$). Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews. These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
This pattern is especially pronounced in categories such as Education and Books & Reference, which show some of the largest gaps between AGR and ANR. In Books & Reference, Gen-AI topics such as AI Performance and Human AI Interaction & Emotional Connection receive high ratings, while non-Gen-AI topics like Technical Difficulties and Customer Support are rated much lower due to crashes, generation failures, and unresolved support issues. Similarly, in Education, users highly rate Gen-AI topics such as Content Quality and Utility & Use Cases, but express dissatisfaction with Monetization Methods & Structure and Technical Difficulties. Together, these results indicate that strong Gen-AI features can substantially boost user satisfaction, but poor reliability, monetization, or support can still pull overall ratings down. Developers should therefore treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
Table 5. Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics (Full dataset considered).
@@
Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category. As shown in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains. Categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively, indicating that users in these domains directly engage with AI features to achieve personal or task-oriented goals such as self-improvement, learning, or workflow assistance.
@@
Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category (Full dataset considered).
In contrast, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews, where 58% of feedback still focuses on non-Gen-AI topics such as usability, technical issues, and subscription or pricing models. This suggests that in creative or media-focused apps, users remain more concerned with traditional performance and value-for-money factors than with AI-driven functionality.
Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting how consistently users discuss Gen-AI features within each domain. Categories such as Entertainment and Health & Fitness exhibit the widest spreads, suggesting substantial variation among apps; some are highly AI-centric, while others receive minimal user attention to Gen-AI features. In contrast, categories like Video Players & Editors and Photography show narrow spreads concentrated at low percentages, indicating that across these domains, Gen-AI features consistently play an assistive or complementary role rather than being the central focus of user experience. Together, these patterns suggest that while certain app categories demonstrate uneven user emphasis on Gen-AI, others achieve relatively uniform user engagement of Gen-AI features.
@@
Observation 2.3: Users discuss a wide range of Gen-AI topics, revealing where developers excel and where improvement is needed. Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively. Developers should pay particular attention to topics that appear most frequently across apps because they represent the aspects users discuss and evaluate most. Excelling in these areas is critical for maintaining competitiveness and user satisfaction. The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality. The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship. Similarly, topics with high average ratings (e.g., Human AI Interaction & Emotional Connection and Utility & use cases) serve as benchmarks for best user experiences that other apps can emulate. Conversely, lower-rated or frequently criticized topics (e.g., Content Policy & Censorship and Content Quality) highlight recurring weaknesses across Gen-AI apps. Developers should treat these as priority areas for refinement to stay ahead of the competition and sustain user trust. We further examine each of the Gen-AI categories in the subsections below.
@@
Table 6. Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
@@
Table 7. Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
@@
However, 33% of reviews reveal concerns in three domains. First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk_Siemenn_Surbehera_Chin_Tyser_Hunter_Raghavan_Hicke_Plummer_Kerret_et_al_2022). Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion. For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour. Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
$\Rightarrow$ Implication 1.
Gen-AI reviews consistently yield higher user ratings across app categories. Developers must recognize Gen-AI as a crucial driver of overall user satisfaction and strategically incorporate these features to remain competitive. However, they should treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing. Even the most advanced AI functionality cannot compensate for poor technical reliability, weak customer support, or frustrating monetization structures. While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts. To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions. Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content. Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
@@
$\Rightarrow$ Implication 3.
Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them. For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences. Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows. They should also add inline editing tools, so users can refine content directly without needing to regenerate responses. Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized. Despite the launch of tools like ChatGPT Edu (openaiChatGPTedu), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features. Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.